A store can sell every gift basket it is able to make. Small baskets contain 5 chocolates, 1 balloon, and 12 roses. Large baskets contain 10 chocolates, 6 balloons, and 12 roses. The shop has access to 80 chocolates, 36 balloons, and 168 roses. The small baskets sell for $20 and the large baskets sells for $40. What combination of baskets should the store make in order to maximize their income? Identify Constraints , objective function, and decision variables

a. Formulate a linear programming model for this problem

b. Solve this method by using graphical analysis.(Excel Method)
Answer #1

Answer:

X1 = number of small basket

X2 = number of large basket

a.

Formulation:

Maximize 20X1 + 40X2

Subject to constraints

5X1 + 10X2 <= 80

X1 + 6X2 <= 36

12X1 + 12X2 <= 168

X1,X2 >= 0

b.

The excel model is shown below

The excel formulas

The parameters

The result